so the issue is complete the square when a does not equal 0.
m^2-m-15=0

Found 3 solutions by josgarithmetic, ikleyn, rothauserc:
Answer by josgarithmetic(39617)
Answer by ikleyn(52778)

.
If you need to complete the square of the left side polynomial, say "PLEASE complete the square of the left side polynomial".

If you need to solve the equation by completing the square, say "PLEASE solve the equation by completing the square".


Use words/terminology correctly.



Answer by rothauserc(4718)

m^2 -m -15 = 0
:
add 15 to both sides of =
:
m^2 -m = 15
:
m^2 -m + (1/4) = 15 + (1/4)
:
(m - (1/2))^2 = 61/4
:
m - (1/2) = square root(61) / 2
:
m = (1/2) + or - square root(61) / 2
:
m = (1 + or - square root(61)) / 2
: